OpenStudy (anonymous):

(-8Sqr40t to the 5th)(5Sqr3t) Help!!!!

OpenStudy (harsimran_hs4):

@inge can you rewrite the expression using equation generator

OpenStudy (anonymous):

\[\left( -8\sqrt{40t ^{5}} \right)\left( 5\sqrt{15t} \right)\]

OpenStudy (harsimran_hs4):

so -8sqrt(40*t^5) * (5sqrt(15t)) rewriting... (-8*5) (sqrt(40) * sqrt(15)) *( sqrt(t^5)*sqrt(t)) clear with the above step ?

OpenStudy (anonymous):

Ok, I get it, each one is placed seperately.

OpenStudy (harsimran_hs4):

ya just regrouping of terms so (8*5) = 40 and sqrt(40)*sqrt(15) = sqrt(40 * 15) = sqrt(600) and sqrt(t^5*t) = sqrt(t^6) = (t^6)^(1/2) = t^(6/2) = t^3 clear?

OpenStudy (anonymous):

Ok I follow...

OpenStudy (harsimran_hs4):

so the expression becomes 40 * sqrt(600) * t^3 sqrt(600) = sqrt(2^3 * 3 * 5^2) = 2*5 sqrt(2*3) = 10sqrt(6) so finally 40* 10sqrt(6)t^3 = 400sqrt(6)t^3

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (harsimran_hs4):

Welcome!!